A study looked at the relationship between alcohol consumption and level of systolic blood pressure (SBP) in women not using oral contraceptives (OC). Alcohol consumption was categorized as follows: no alcohol use; ≤ 10 oz/week alcohol consumption; > 10 oz/week alcohol consumption. The results for the SBP measurements for women 30-39 years of age are given below. Group Mean var n A. No Alcohol Use 105.3 10.6 24 B. ≤ 10 oz/week alcohol consumption 110.4 13.4 12 C. > 10 oz/week alcohol consumption 122.5 12.5 10 Conduct a two sample independent t-test to determine if women OC users ages 30-39 that consume ≤ 10 oz/week of alcohol (define as Group 1) have a significantly higher SBP than women who consumed no alcohol (define as Group 2) with an alpha level of 0.05. The pooled standard deviation s is 3.39. What is the critical value? Round to 4 decimal places. a. t=1.6909 b. t=1.96 c. t=1.645 d. t=1.6939
Answer #1

Degrees of freedom = n1 + n2 - 2 = 24 + 12 - 2 = 34

Hence,

Critical t value = 1.6909

Option A is correct.